# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v1

@@ A Service Driver Based Application Execution and Development Method in Multi-cloud Context
Monitoring-Aware Optimal Deployment for Applications Based on Microservices
IEEE Transactions on Services Computing
2021 (Fadda et al., 2019)
S47
@@
MSSF: A Step towards User-Friendly Multi-cloud Data Dispersal
International Conference on Cloud Computing (CLOUD)
2014 (Libardi et al., 2014)
S48
@@
Multi-agent QoS-aware autonomic resource provisioning framework for elastic BPM in containerized multi-cloud environment
Journal of Ambient Intelligence and Humanized Computing
2022 (Saif et al., 2023)
S49
@@
Multicloud- auto scale with prediction and delta correction algorithm
International Conference on Intelligent Computing, Instrumentation and Control Technologies (ICICICT)
2019 (Harwalkar et al., 2019b)
50
@@
Multi-cloud Distribution of Virtual Functions and Dynamic Service Deployment: Open ADN Perspective
IEEE International Conference on Cloud Engineering (IC2E)
2015 (Bhamare et al., 2015)
S51
@@
Multi-cloud resource provisioning with Aneka: A unified and integrated utilization of Microsoft Azure and Amazon EC2
International Conference on Computing and Network Communications (CoCoNet)
2015 (Buyya and Barreto, 2015)
S52
MUSA Deployer: Deployment of Multi-cloud Applications
@@
IEEE International Workshop on Enabling Technologies: Infrastructure for Collaborative Enterprises (WETICE)
2017 (Casola et al., 2017)
S53
@@
NoMISHAP: A Novel Middleware Support for High Availability in Multicloud PaaS
IEEE Cloud Computing
2017 (Acquaviva et al., 2017)
S54
Object storage in the cloud and multi-cloud: State of the art and the research challenges
@@
IEEE International Conference on Automation, Quality and Testing, Robotics (AQTR)
2018 (Bucur et al., 2018)
S55
On the Deployment of an Open-Source, 5G-Aware Evaluation Testbed
@@
International Conference on Mobile Cloud Computing, Services, and Engineering (MobileCloud)
2018 (Bolivar et al., 2018)
S56
OpenStack Federation in Experimentation Multi-cloud Testbeds
@@
IEEE International Conference on Cloud Computing Technology and Science (CloudCom)
2013 (del Castillo et al., 2013)
S57
@@
Orchestrated Multi-Cloud Application Deployment in OpenStack with TOSCA
International Conference on Smart Computing (SMARTCOMP)
2017 (Tricomi et al., 2017)
S58
Orchestration Based Hybrid or Multi Clouds and Interoperability Standardization
@@
IEEE International Conference on Cloud Computing in Emerging Markets (CCEM)
2018 (Sitaram et al., 2018)
S59
@@
Performance Analysis of Zero-Trust multi-cloud
IEEE International Conference on Cloud Computing (CLOUD)
2021 (Rodigari et al., 2021)
S60
@@
Policies Based Container Migration Using Cross-Cloud Management Platform
International Conference on Information and Automation for Sustainability (ICIAfS)
2018 (Janarthanan et al., 2018)
S61
@@
Policy-based deployment in a hybrid and multi-cloud environment
International Conference on Cloud Computing and Services Science (CLOSER)
2019 (Di Modica et al., 2019)
S62
@@
Predicting the End-to-End Tail Latency of Containerized Microservices in the Cloud
International Conference on Cloud Engineering (IC2E)
2019 (Rahman and Lama, 2019)
S63
@@
Preventing vendor lock-ins via an interoperable multi-cloud deployment approach
International Conference for Internet Technology and Secured Transactions (ICITST)
2017 (Pellegrini et al., 2017)
S64
@@
Private STaaS with OpenStack Cinder Volumes for Hybrid/Multi-cloud
International Conference on Advances in Computing and Communication Engineering (ICACCE)
2019 (Harwalkar et al., 2019a)
S65
@@
Real-World Experience with a Multicloud Exchange
IEEE Cloud Computing
2017 (Zeck and Bouroudjian, 2017)
S66
@@
Reliability Management for Blockchain-Based Decentralized Multi-Cloud
International Symposium on Cluster Computing and the Grid (CCGRID)
2020 (Aral et al., 2020)
S67
@@
Reliability-Aware Cost-Efficient Scientific Workflows Scheduling Strategy on Multi-Cloud Systems
IEEE Transactions on Cloud Computing
2022 (Tang, 2021)
S68
@@
Resource optimization of container orchestration: a case study in multi-cloud microservices-based applications
The Journal of Supercomputing
2018 (Guerrero et al., 2018b)
S69
@@
Scalable multi-cloud platform to support industry and scientific applications
International Convention on Information and Communication Technology, Electronics and Microelectronics (MIPRO)
2018 (Kiss, 2018)
S70
@@
Search-based Methods for Multi-Cloud Configuration
International Conference on Cloud Computing (CLOUD)
2022 (Lazuka et al., 2022)
S71
@@
Secure access control for multi-cloud resources
IEEE Conference on ALocal Computer Networks Workshops (LCN Workshops)
2015 (Graupner et al., 2015b)
S72
Secure and Dependable Multi-Cloud Network Virtualization
@@
International Workshop on Security and Dependability of Multi-Domain Infrastructures
2017 (Alaluna et al., 2017)
S73
@@
Secure and Parallel Expressive Search over Encrypted Data with Access Control in Multi-CloudIoT
Cloudification of the Internet of Things (CIoT)
2018 (Guechi and Maamri, [n. d.])
S74
@@
Self-healing cloud services in private multi-clouds
International Conference on High Performance Computing & Simulation (HPCS)
2018 (Mfula and Nurminen, 2018)
S75
@@
SpyStorage: A Highly Reliable Multi-Cloud Storage with Secure and Anonymous Data Sharing
International Conference on Networking, Architecture, and Storage (NAS)
2017 (Shen et al., 2017)
S76
@@
Support Services for Applications Execution in Multi-clouds Environments
International Conference on Autonomic Computing (ICAC)
2016 (Pop et al., 2016)
S77
SynAPTIC: Secure and persistent connectivity for containers
@@
International Symposium on Cluster, Cloud and Grid Computing (CCGRID)
2017 (Ranjbar et al., 2017)
S78
@@
System Restore in a Multi-cloud Data Pipeline Platform
International Conference on Dependable Systems and Networks – Industry Track
2019 (Wang et al., 2019)
S79
@@
Systematic and Recomputable Comparison of Multi-cloud Management Platforms
International Conference on Cloud Computing Technology and Science (CloudCom)
2018 (Serhiienko and Spillner, 2018)
S80
@@
Towards a lightweight multi-cloud DSL for elastic and transferable cloud-native applications
International Conference on Cloud Computing and Services Science (CLOSER)
2018 (Quint and Kratzke, 2018)
S81
@@
Towards Formal-Based Semantic Interoperability in Multi-Clouds: The FCLOUDS Framework
IEEE International Conference on Cloud Computing (CLOUD)
2017 (Challita et al., 2017)
S82
@@
Towards Model-Driven Provisioning, Deployment, Monitoring, and Adaptation of Multi-cloud Systems
IEEE International Conference on Cloud Computing (CLOUD)
2013 (Ferry et al., 2013)
S83
@@
Towards Trustworthy Multi-Cloud Services Communities: A Trust-Based Hedonic Coalitional Game
IEEE Transactions on Services Computing
2018 (Wahab et al., 2016)
S84
@@
Multi-cloud provisioning of business processes
Journal of Cloud Computing
2019 (Kritikos et al., 2019)
S85
@@
Evaluating the DevOps Reference Architecture for Multi-cloud IoT-Applications
SN Computer Science
2021 (Bou Ghantous and Gill, 2021)
S86
@@
Hybrid IT and Multi Cloud an Emerging Trend and Improved Performance in Cloud Computing
SN Computer Science
2020 (Gundu et al., 2020)

### v2

@@ 8. Appendices
S46 Monitoring-Aware Optimal Deployment for Applications Based on Microservices IEEE Transactions on Services Computing 2021 (Fadda et al., 2019)
@@
S47 MSSF: A Step towards User-Friendly Multi-cloud Data Dispersal International Conference on Cloud Computing (CLOUD) 2014 (Libardi et al., 2014)
@@
S48 Multi-agent QoS-aware autonomic resource provisioning framework for elastic BPM in containerized multi-cloud environment Journal of Ambient Intelligence and Humanized Computing 2022 (Saif et al., 2023)
@@
S49 Multicloud- auto scale with prediction and delta correction algorithm International Conference on Intelligent Computing, Instrumentation and Control Technologies (ICICICT) 2019 (Harwalkar et al., 2019b)
@@
50 Multi-cloud Distribution of Virtual Functions and Dynamic Service Deployment: Open ADN Perspective IEEE International Conference on Cloud Engineering (IC2E) 2015 (Bhamare et al., 2015)
@@
S51 Multi-cloud resource provisioning with Aneka: A unified and integrated utilization of Microsoft Azure and Amazon EC2International Conference on Computing and Network Communications (CoCoNet) 2015 (Buyya and Barreto, 2015)
@@
S52 MUSA Deployer: Deployment of Multi-cloud Applications IEEE International Workshop on Enabling Technologies: Infrastructure for Collaborative Enterprises (WETICE) 2017 (Casola et al., 2017)
@@
S53 NoMISHAP: A Novel Middleware Support for High Availability in Multicloud PaaS IEEE Cloud Computing 2017 (Acquaviva et al., 2017)
@@
S54 Object storage in the cloud and multi-cloud: State of the art and the research challengesIEEE International Conference on Automation, Quality and Testing, Robotics (AQTR) 2018 (Bucur et al., 2018)
@@
S55 On the Deployment of an Open-Source, 5G-Aware Evaluation Testbed International Conference on Mobile Cloud Computing, Services, and Engineering (MobileCloud) 2018 (Bolivar et al., 2018)
@@
S56 OpenStack Federation in Experimentation Multi-cloud Testbeds IEEE International Conference on Cloud Computing Technology and Science (CloudCom) 2013 (del Castillo et al., 2013)
@@
S57 Orchestrated Multi-Cloud Application Deployment in OpenStack with TOSCA International Conference on Smart Computing (SMARTCOMP) 2017 (Tricomi et al., 2017)
@@
S58 Orchestration Based Hybrid or Multi Clouds and Interoperability Standardization IEEE International Conference on Cloud Computing in Emerging Markets (CCEM) 2018 (Sitaram et al., 2018)
@@
S59 Performance Analysis of Zero-Trust multi-cloud IEEE International Conference on Cloud Computing (CLOUD) 2021 (Rodigari et al., 2021)
@@
S60 Policies Based Container Migration Using Cross-Cloud Management Platform International Conference on Information and Automation for Sustainability (ICIAfS) 2018 (Janarthanan et al., 2018)
@@
S61 Policy-based deployment in a hybrid and multi-cloud environment International Conference on Cloud Computing and Services Science (CLOSER) 2019 (Di Modica et al., 2019)
@@
S62 Predicting the End-to-End Tail Latency of Containerized Microservices in the Cloud International Conference on Cloud Engineering (IC2E) 2019 (Rahman and Lama, 2019)
@@
S63 Preventing vendor lock-ins via an interoperable multi-cloud deployment approach International Conference for Internet Technology and Secured Transactions (ICITST) 2017 (Pellegrini et al., 2017)
@@
S64 Private STaaS with OpenStack Cinder Volumes for Hybrid/Multi-cloud International Conference on Advances in Computing and Communication Engineering (ICACCE) 2019 (Harwalkar et al., 2019a)
@@
S65 Real-World Experience with a Multicloud Exchange IEEE Cloud Computing 2017 (Zeck and Bouroudjian, 2017)
@@
S66 Reliability Management for Blockchain-Based Decentralized Multi-Cloud International Symposium on Cluster Computing and the Grid (CCGRID) 2020 (Aral et al., 2020)
@@
S67 Reliability-Aware Cost-Efficient Scientific Workflows Scheduling Strategy on Multi-Cloud Systems IEEE Transactions on Cloud Computing 2022 (Tang, 2021)
@@
S68 Resource optimization of container orchestration: a case study in multi-cloud microservices-based applicationsThe Journal of Supercomputing 2018 (Guerrero et al., 2018a)
@@
S69 Scalable multi-cloud platform to support industry and scientific applications International Convention on Information and Communication Technology, Electronics and Microelectronics (MIPRO) 2018 (Kiss, 2018)
@@
S70 Search-based Methods for Multi-Cloud Configuration International Conference on Cloud Computing (CLOUD) 2022 (Lazuka et al., 2022)
@@
S71 Secure access control for multi-cloud resources IEEE Conference on ALocal Computer Networks Workshops (LCN Workshops) 2015 (Graupner et al., 2015a)
@@
S72 Secure and Dependable Multi-Cloud Network Virtualization International Workshop on Security and Dependability of Multi-Domain Infrastructures 2017 (Alaluna et al., 2017)
@@
S73 Secure and Parallel Expressive Search over Encrypted Data with Access Control in Multi-CloudIoT Cloudification of the Internet of Things (CIoT) 2018 (Guechi and Maamri, [n. d.])
@@
S74 Self-healing cloud services in private multi-clouds International Conference on High Performance Computing & Simulation (HPCS) 2018 (Mfula and Nurminen, 2018)
@@
S75 SpyStorage: A Highly Reliable Multi-Cloud Storage with Secure and Anonymous Data Sharing International Conference on Networking, Architecture, and Storage (NAS) 2017 (Shen et al., 2017)
@@
S76 Support Services for Applications Execution in Multi-clouds Environments International Conference on Autonomic Computing (ICAC) 2016 (Pop et al., 2016)
@@
S77 SynAPTIC: Secure and persistent connectivity for containers International Symposium on Cluster, Cloud and Grid Computing (CCGRID) 2017 (Ranjbar et al., 2017)
@@
S78 System Restore in a Multi-cloud Data Pipeline Platform International Conference on Dependable Systems and Networks – Industry Track 2019 (Wang et al., 2019)
@@
S79 Systematic and Recomputable Comparison of Multi-cloud Management Platforms International Conference on Cloud Computing Technology and Science (CloudCom) 2018 (Serhiienko and Spillner, 2018)
@@
S80 Towards a lightweight multi-cloud DSL for elastic and transferable cloud-native applications International Conference on Cloud Computing and Services Science (CLOSER) 2018 (Quint and Kratzke, 2018)
@@
S81 Towards Formal-Based Semantic Interoperability in Multi-Clouds: The FCLOUDS Framework IEEE International Conference on Cloud Computing (CLOUD) 2017 (Challita et al., 2017)
@@
S82 Towards Model-Driven Provisioning, Deployment, Monitoring, and Adaptation of Multi-cloud Systems IEEE International Conference on Cloud Computing (CLOUD) 2013 (Ferry et al., 2013)
@@
S83 Towards Trustworthy Multi-Cloud Services Communities: A Trust-Based Hedonic Coalitional Game IEEE Transactions on Services Computing 2018 (Wahab et al., 2016)
@@
S84 Multi-cloud provisioning of business processes Journal of Cloud Computing 2019 (Kritikos et al., 2019)
@@
S85 Evaluating the DevOps Reference Architecture for Multi-cloud IoT-Applications SN Computer Science 2021 (Bou Ghantous and Gill, 2021)
@@
S86 Hybrid IT and Multi Cloud an Emerging Trend and Improved Performance in Cloud Computing SN Computer Science 2020 (Gundu et al., 2020)